OpenStudy (anonymous):

What is the common difference of a 57-term arithmetic sequence where the first term is -25 and the sum is 17,727?

OpenStudy (anonymous):

first we need a formula to show the sum... \[S = \frac{ n(t _{1} + t _{n}) }{ 2 }\]

OpenStudy (anonymous):

which "n" is the number of the terms , "t1" = the first term and "tn" = the last (or the n term)

OpenStudy (anonymous):

So I got 647 for n...what next? @PFEH.1999

OpenStudy (anonymous):

how did u find tn?

OpenStudy (anonymous):

i meant tn...that's 647

OpenStudy (anonymous):

n was 57

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

ok let me check it.

OpenStudy (anonymous):

yess correct , tn = 47 and \( t _{n} = t _{1} + (n-1) d\) and d is the common difference

OpenStudy (anonymous):

@ambiguousmathstudent have you want d or not?

OpenStudy (anonymous):

I got a weird answer, i think i did it wrong. i have 47=-25+(57-1)d and i got a weird answer when i solved for d

OpenStudy (anonymous):

the answer should be one of these: 11 11.5 12 12.5

OpenStudy (anonymous):

47??? tn = 647 !

OpenStudy (anonymous):

hahahhaa my bad...i have awful handwriting. let me try that again

OpenStudy (anonymous):

wait i still got -2.16 which isnt an answer @PFEH.1999

OpenStudy (anonymous):

647 + 25 = 672 = 56d so d = 672/56b = 12 ;)

OpenStudy (anonymous):

you've made a mistake ;)

OpenStudy (anonymous):

ohhhhh i see what i did wrong, thank you so much. you're the best ((:

OpenStudy (anonymous):

you're welcome ;) thank u ;) :)